3. The cash account of JM Company has a beginning balance of P129,937.50. Its year-end balance stands at P254,925. The table below shows the summarized transactions from the cash account of JM Company.

Cash received from customers P 725,175.00

Payments to suppliers 300,547.50

Payments for other operating expenses 58,850.00

Salaries paid 19,992.50

Purchase of equipment 253,000.00

Sale of delivery equipment 6,875.00

Proceeds from sale of computer equipment 47,762.50

Withdrawals of owners 229,350.00

Proceeds from bank loan 82,500.00

Contributions of owners 137.500.00

Interest paid 8,085.00

Requirements:

1. How much is cash flow from operating activities?

2. How much is cash flow from investing activities?

3. How much is cash flow from financing activities?

4. Is JM Company’s statement of cash flows reconciles with the balance of the cash account as of the end of the year? Show your computation.



1
Expert's answer
2021-03-03T11:07:33-0500

1.=725,175+(-300,547.50)+(-58,850)+(-19,992.50)

=345,785


2.=-253,000+6,875.00+47,762.50

=-198,362.5


3.=82,500+137,500+(-229,350)+(-8,085.00)

=-17,435


4.The two do not reconcile


Bal at the end of yr = 254,925.00

Bal at the beginning of yr =129,937.50


Total Cash flows=345,785 +(-198,362.5)+(-17,435.00)

= 129,987.50

Bal at end of yr =129,937.50+129,987.50

=259,925.00


254,925.00 is not equal to 259,925